max buys 7 plam trees. He buys 10 more cactus plants than he buys palm trees. He buys 12 more ferns than he buys cactuses. Write a number sentence to shoe how many ferns max buys?
Answer by John10(297)

Max buys 7 palm trees. He buys 10 more cactus plants than he buys palm trees. He buys 12 more ferns than he buys cactus. Write a number sentence to shoW how many ferns max buys?
Suppose the number of palm trees are x then x = 7
Max buys 10 more cactus plants than he buys palm trees: x + 10 = 7 + 10 = 17 cactus.
He buys 12 more ferns than he buys cactus: 17 + 12 = 29 ferns
PLEASE MAKE SURE TO CHECK WHETHER THIS IS THE WAY THAT YOUR PROF. REQUIRES!